# GlaGAN: A Generative Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

### v4

@@ 4.1.1 GSS-RetVein
GSS-RetVein OCTA-500 ROSE
@@
U-Net 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
U-Net++ 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
Attention U-Net 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
Mask R-CNN 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
YOLOv11-x 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
MedSAM 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
CauSSL 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
UniverSeg 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
S2VNet 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
Tyche 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
Table 2: Conducting Mixed Validation on Different Models using Mixed Datasets ($\times$100%)
@@
ROSE GSS-RetVein OCTA-500 GSS-RetVein (Test) OCTA-500 (Test)
GSS-RetVein OCTA-500 ROSE OCTA-500 (Train) GSS-RetVein (Train)
3mm 6mm 3mm 6mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
U-Net 98.34 99.16 98.32 99.15 93.89 96.85 94.23 97.03 96.21 98.07 95.23 97.56 96.98 98.47 96.81 98.38
U-Net++ 98.99 99.49 98.78 99.39 94.08 96.95 94.46 97.15 95.73 97.82 97.01 98.51 97.59 98.78 96.93 98.44
Attention U-Net 98.54 99.26 98.23 99.11 94.07 96.94 94.52 97.18 96.99 98.47 96.09 98.01 97.21 98.59 96.78 98.36
Mask R-CNN 96.37 98.15 96.31 98.12 90.73 95.14 91.89 95.77 94.23 97.03 94.11 96.97 96.68 98.31 95.96 97.94
YOLOv11-x 98.88 99.44 98.72 99.36 94.11 96.97 94.49 97.17 97.32 98.64 96.54 98.24 96.12 98.02 96.63 98.29
MedSAM 99.04 99.52 98.98 99.49 95.01 97.41 95.42 97.66 97.54 98.75 95.87 97.89 98.23 99.11 97.46 98.72
CauSSL 96.83 98.38 95.89 97.90 96.98 98.47 97.83 98.88 99.12 99.56 98.81 99.40 94.87 97.33 95.34 97.62
UniverSeg 98.96 99.48 98.56 99.28 94.23 97.03 94.59 97.24 96.81 98.38 96.32 98.16 97.91 98.96 98.12 99.05
S2VNet 99.23 99.61 98.79 99.39 93.89 96.85 95.22 97.55 96.95 98.44 95.83 97.87 97.87 98.95 97.87 98.95
Tyche 98.78 99.38 98.43 99.21 94.22 97.02 94.98 97.46 96.88 98.41 96.11 98.02 97.84 98.94 97.68 98.87
GlaGAN 99.55 99.78 99.21 99.60 96.98 98.46 97.69 98.83 98.87 99.43 98.79 99.39 99.58 99.79 99.41 99.70
Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
4.1.2 OCTA-500
OCTA-500 [1] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies. It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
4.1.3 ROSE
ROSE [4] is an open-source collection designed for retinal blood vessel segmentation using OCTA images. It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations. ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
4.2 Implement
Table 3: Comparison Vessel Structure Refinement Images with Real OCTA Images ($\times$100%)
Model Dataset
GSS-RetVein OCTA-500 ROSE
@@ 4.2 Implement
GAN 91.76 5.69 90.39 4.43 89.69 9.85 91.06 7.36 87.84 8.73
CycleGAN 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
GAN 93.62 3.29 93.87 3.79 91.28 6.98 93.19 5.54 91.73 4.18
CycleGAN 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
4.2.1 Hyperparameter
for GlaGAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. GlaGAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
4.2.2 Hardware
the hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
4.2.3 Evaluation Metric
Table 4: Ablation Experiment of GlaGAN ($\times$100%)
@@ 4.2.3 Evaluation Metric
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
GAN U-Net 92.89 96.31 92.44 96.07 93.31 96.56 92.84 96.31 93.37 96.59
GAN U-Net++ 93.62 96.70 92.87 96.30 93.72 96.73 93.03 96.36 92.84 96.31
GAN Attention U-Net 93.56 96.67 93.12 96.44 93.65 96.70 92.85 96.31 93.42 96.61
GAN MedSAM 93.73 96.76 92.76 96.24 92.83 96.30 92.47 96.10 93.76 96.75
GAN DFS 95.81 97.86 95.24 97.56 94.83 97.33 94.87 97.35 94.73 97.27
CycleGAN U-Net 93.44 96.61 93.42 96.60 94.73 97.26 93.97 96.89 93.76 96.75
CycleGAN U-Net++ 93.72 96.76 93.62 96.70 94.58 97.18 94.36 97.05 93.19 96.46
CycleGAN Attention U-Net 93.79 96.80 93.45 96.61 94.59 97.18 94.29 97.01 94.03 96.89
CycleGAN MedSAM 94.02 96.92 94.37 97.10 93.83 96.72 94.37 97.19 94.18 96.96
CycleGAN DFS 96.23 98.08 96.42 98.18 96.68 98.33 96.27 98.13 95.83 97.84
CycleGAN U-Net 94.56 97.20 94.52 97.18 94.86 97.31 94.79 97.27 94.19 96.94
CycleGAN U-Net++ 94.98 97.43 94.36 97.10 94.83 97.29 95.09 97.41 94.27 96.98
CycleGAN Attention U-Net 95.06 97.47 95.12 97.50 95.07 97.39 94.97 97.33 94.39 97.04
CycleGAN MedSAM 95.33 97.61 95.21 97.55 95.36 97.55 95.11 97.41 94.82 97.27
GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
5 Experimental Result
5.1 Experimental Setup
We selected many SOTA models for experimental verification, which are: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15], Tyche [16], GAN [34] and CycleGAN [35]. Among them, we used tool Optuna [^4] https://optuna.org/ to optimize their parameters. All final experimental results are the average of ten cross-validation times. * means models’ hyperparameters have been optimized by us.
5.2 Comparative Experiment
As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, GlaGAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation. As shown in Table 2, the second row is the test data set, and the third row is the training data set. Models trained on GSS-RetVein and OCTA-500 were tested on ROSE, where GlaGAN achieved the highest IoU and Dice scores, outperforming S2VNet, CauSSL, and MedSAM. For generalization, training on ROSE and testing on GSS-RetVein and OCTA-500 confirmed GlaGAN’s superiority, followed by MedSAM and CauSSL. Across different scanning resolutions, GlaGAN consistently led in segmentation accuracy, with MedSAM and UniverSeg performing well, while U-Net variants lagged. These results affirm GSS-RetVein’s quality and GlaGAN’s segmentation excellence.
5.3 Ablation Experiment
For baseline GAN model, GAN and CycleGAN are selected. For baseline segmentor, U-Net, U-Net++, Attention U-Net and MedSAM are selected. For GlaGAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Table 4, since main blood vessels make up a small proportion, the value calculated is also relatively low. When the minimum radius ratio is found, like GlaGAN, aligning with biostatistical properties [29, 30, 31, 33, 32], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Table 5: Ablation Experiment of Hyperparameters ($\times$100%)
Hyperparameter Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
Rmin 0.00 train 1.03 2.04 1.00 1.98 1.31 2.59 1.22 2.41 1.13 2.23
test 0.75 1.49 0.87 1.73 0.91 1.80 0.77 1.52 0.93 1.84
0.10 train 41.32 58.47 41.98 59.12 43.09 60.23 45.57 62.59 44.29 61.38
test 40.29 57.42 42.75 59.89 42.31 59.45 45.06 62.13 43.97 61.08
0.20 train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
0.30 train 87.15 93.13 86.29 92.64 86.04 92.49 85.74 92.30 84.37 91.52
test 84.17 91.40 83.96 91.27 84.09 91.35 84.02 91.31 83.98 91.28
0.40 train 57.92 73.34 56.44 72.15 55.86 71.65 55.45 71.28 57.09 72.69
test 58.73 73.98 57.85 73.29 57.29 72.85 56.93 72.53 57.24 72.80
0.50 train 55.72 71.56 56.98 72.60 54.79 70.78 56.93 72.53 57.04 72.65
test 54.17 70.25 53.83 69.99 52.62 68.97 53.71 69.89 53.83 69.99
0.60 train 40.26 57.40 41.27 58.42 40.39 57.53 41.03 58.17 39.27 56.39
test 35.62 52.51 38.92 56.03 37.46 54.50 37.08 54.10 35.74 52.63
0.70 train 35.36 52.25 34.72 51.55 33.91 50.64 37.06 54.08 35.94 52.85
test 32.24 48.76 31.83 48.27 32.64 49.21 33.39 50.06 34.07 50.83
0.80 train 17.83 30.26 19.25 32.29 18.83 31.69 17.79 30.21 16.68 28.59
test 26.37 41.73 25.73 40.93 25.49 40.61 25.13 40.15 24.80 39.73
0.90 train 9.85 17.94 10.32 18.71 9.89 17.99 11.27 20.26 9.91 18.03
test 12.24 21.82 13.06 23.10 12.94 22.92 11.98 21.39 10.57 19.12
1.00 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Node 0 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
5000 train 6.34 11.92 7.81 14.49 6.28 11.82 6.43 12.09 7.02 13.12
test 6.06 11.43 5.86 11.07 5.93 11.19 6.18 11.64 6.58 12.35
10000 train 17.82 30.24 18.92 31.81 17.96 30.44 19.06 32.02 17.48 29.75
test 20.59 34.15 21.76 35.73 22.84 37.19 21.96 36.00 20.98 34.69
15000 train 20.58 34.14 21.96 36.00 20.84 34.50 21.37 35.21 20.71 34.29
test 28.73 44.63 29.64 45.71 30.74 47.04 28.93 44.87 29.74 45.83
20000 train 41.47 58.61 42.34 59.49 41.65 58.80 42.07 59.18 44.03 61.14
test 42.06 59.17 41.11 58.26 41.93 59.04 40.62 57.77 41.98 59.12
25000 train 49.38 66.12 52.14 68.53 50.39 66.96 52.97 69.26 51.07 67.60
test 52.01 68.41 51.49 67.98 51.76 68.21 52.44 68.79 53.06 69.35
30000 train 59.23 74.39 59.17 74.34 58.48 73.78 57.99 73.40 58.49 73.79
test 61.25 75.97 60.84 75.64 59.73 74.78 61.79 76.37 60.83 75.63
35000 train 62.33 76.79 63.34 77.56 62.15 76.64 61.96 76.49 62.15 76.64
test 62.51 76.96 64.08 78.12 63.24 77.47 62.44 76.89 63.19 77.43
40000 train 62.74 77.13 64.27 78.25 63.74 77.88 63.96 78.06 63.49 77.64
test 62.59 77.03 64.13 78.16 63.96 78.06 61.84 76.39 63.97 78.07
45000 train 63.37 77.57 65.03 78.81 65.17 79.57 65.03 78.81 65.93 79.41
test 65.72 79.30 66.34 79.77 64.44 78.39 65.39 78.95 66.67 80.00
50000 train 66.53 79.89 66.02 79.53 65.52 79.14 65.74 79.31 66.03 79.54
test 69.34 81.89 68.18 81.06 69.01 81.65 66.95 80.19 68.39 81.24
50000 train 71.24 83.20 72.47 84.02 69.08 81.71 68.79 81.52 69.28 81.83
test 73.69 84.85 72.61 84.12 74.17 85.15 75.36 85.94 74.13 85.10
55000 train 75.48 85.99 76.24 86.54 76.14 86.47 75.83 86.25 74.19 85.17
test 76.02 86.38 75.47 85.98 74.45 85.34 73.99 85.05 74.73 85.57
60000 train 79.28 88.47 78.96 88.23 79.48 88.61 77.20 87.13 75.96 86.35
test 78.49 87.94 77.85 87.52 77.02 87.03 79.36 88.52 75.89 86.29
65000 train 85.27 92.05 86.09 92.52 85.78 92.29 85.96 92.42 86.06 92.50
test 84.18 91.40 84.07 91.33 83.33 90.89 82.39 90.37 83.09 90.76
70000 train 97.46 98.71 97.28 98.62 97.08 98.52 96.96 98.46 97.07 98.51
test 98.32 99.15 98.29 99.14 99.03 99.51 99.01 99.50 98.76 99.37
…… train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
As shown in Table 5, the optimal choice of $R_{min}$ (0.2) enables GlaGAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%. When $R_{min}=1$, GlaGAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold. When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
Also, as shown in Fig. ref1 and Table 5, a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of GlaGAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar. Consequently, their curves largely overlap, with GlaGAN exhibiting a slight increase.
5.4 Evaluation of Vessel Structure Refinement
@@ 5.5 Efficiency Evaluation
As shown in Table 6, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
Table 6: Comparison of Performance Indicators between DFS and SLSMs ($\times$100%)
Indicator SLSMs DFS (ours)
GPU $\checkmark$ $\times$
Annotation $\checkmark$ $\times$
Training $\checkmark$ $\times$
Explainability $\times$ $\checkmark$
2D and 3D $\times$ $\checkmark$
Generalization $\times$ -
@@
U-Net 21ms 31M 98 low
U-Net++ 62ms 76M 304 high
Attention U-Net 83ms 62M 215 slightly high
MedSAM 425ms 109M 1250 extremely high
DFS (CPU) 0.018ms 0M 0.00028 -
DFS (GPU) $\leq$ 0.01ms 0M 0.00028 extremely low
Table 7: Comparison of Performance Details between DFS and SLSMs
@@ 5.6 Visualization
As shown in Fig. 6, 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 6(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 6(e) preserving vessel integrity and offering a clearer structure than the original image (Fig. 6(c)).
@@
(b) Original Image (2D)
(c) Original Image (3D)
(d) Segmentation (2D)
(e) Segmentation (3D)
Figure 6: Segmentation Results of GlaGAN on GSS-RetVein (2D and 3D)

### v5

@@ 5.1 Comparative Experiment
UTSW OCTA-500 ROSE
@@
U-Net [10] 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
U-Net++ [7] 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
Attention U-Net [12] 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
Mask R-CNN [9] 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
YOLOv11-x [11] 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
MedSAM [8] 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
CauSSL [15] 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
UniverSeg [16] 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
S2VNet [17] 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
Tyche [19] 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, X-GAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
Table 2: Comparison Vessel Structure Refinement with Real OCTA ($\times$100%)
@@
UTSW OCTA-500 ROSE
@@
CycleGAN [39] 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
Human Evaluation: the medical team verified the corrected images. Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
5.2 Ablation Eperiment
For baseline GAN model, GAN [38] and CycleGAN [39] are selected. For baseline segmentor, U-Net [10], U-Net++ [7], Attention U-Net [12] and MedSAM [8] are selected. For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Table 3, since main blood vessels make up a small proportion, the value calculated is also relatively low, aligning with biostatistical properties [34, 35, 36, 37], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Table 3: Ablation Experiment of X-GAN ($\times$100%)
@@ 5.2 Ablation Eperiment
UTSW
6mm 3mm
IoU Dice IoU Dice
GAN [38] + U-Net [10] 92.89 96.31 92.44 96.07
GAN [38] + U-Net++ [7] 93.62 96.70 92.87 96.30
GAN [38] + Attention U-Net [12] 93.56 96.67 93.12 96.44
GAN [38] + MedSAM [8] 93.73 96.76 92.76 96.24
GAN [38] + DFS ($R_{min}$=0.2) 95.81 97.86 95.24 97.56
CycleGAN [39] + U-Net [10] 93.44 96.61 93.42 96.60
CycleGAN [39] + U-Net++ [7] 93.72 96.76 93.62 96.70
CycleGAN [39] + Attention U-Net [12] 93.79 96.80 93.45 96.61
CycleGAN [39] + MedSAM [8] 94.02 96.92 94.37 97.10
CycleGAN [39] + DFS ($R_{min}$=0.2) 96.23 98.08 96.42 98.18
CycleGAN (ours) + U-Net [10] 94.56 97.20 94.52 97.18
CycleGAN (ours) + U-Net++ [7] 94.98 97.43 94.36 97.10
CycleGAN (ours) + Attention U-Net [12] 95.06 97.47 95.12 97.50
CycleGAN (ours) + MedSAM [8] 95.33 97.61 95.21 97.55
X-GAN 99.41 99.71 98.66 99.33
As shown in Fig. 3(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%. When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold. When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
(a) $R_{min}$
(b) The Number of Nodes NN
Figure 3: Ablation Experiments of Key Parameters ($\times$100%)
Also, as shown in Fig. 3(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of X-GAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
@@ 5.3 Efficiency Evaluation
Compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
Table 4: Comparison of Performance Details between DFS (our) and SLSMs
@@
U-Net [10] 21ms 31M 98 low
U-Net++ [7] 62ms 76M 304 high
Attention U-Net [12] 83ms 62M 215 slightly high
MedSAM [8] 425ms 109M 1250 extremely high
DFS (ours) 0.018ms 0M 0.00028 -
DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
@@ 5.4 Visualization
As shown in Fig. 4, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 4(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 4(c) preserving vessel integrity and offering a clearer structure than the original image (Fig. 4(b)).
@@
(b) Original Image (3D)
(c) Segmentation (3D)
Figure 4: Segmentation of X-GAN (2D and 3D)